OpenStudy (anonymous):

Find the center and radius of x2 + y2 + 2x - 10y + 10 = 0

OpenStudy (anonymous):

factorise it to get (x+1)^2 + (y-5)^2 = 16 = 4^2 hence centre is (-1,5) and radius is 4

OpenStudy (anonymous):

centre(-1,5) radius=4 using d general eqn of a circle=x^2+y^2-2ax-2bx+c=0 wen d centre is(a,b) radius=sqrt a^2+b^2-c

OpenStudy (anonymous):

x^2+y^2+2x-10y+10=0 ==>(x+1)^2+(y-5)^2+10-1-25=0 ==>(x+1)^2+(y-5)^2=4^2 ==>center is (-1,5), and rad=4